When a sample of  C2H5OH  was combusted, the volume of  CO2(g)  produced was 18.0L when measured at 21.7°C and 1.03atm . Determine the number of moles of  CO2(g)  that was produced.
1
Expert's answer
2020-11-11T10:48:50-0500

Assume ideal gas and use the law:

p*V = n*R*T

p = 1.03*101.325 = 104.36475 kPa

V = 18.0 L

R = 8.314 J/(K*mole)

T = 21.7+273.15 = 294.85 K

n = 18*104.36475/(8.314*300.25) = 0.77 mole

Number of CO2(g) moles is 0.77